OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6 -14 24 14 -24

OpenStudy (anonymous):

Manipulate the equations so that y is the only variable: 3x - 6y = 12 -2x + 3y = 6 Then: -2x = 6 - 3y x = (6 - 3y)/(-2)

OpenStudy (anonymous):

And then substitute x into the first equation: 3((6 - 3y)/(-2)) - 6y = 12

OpenStudy (anonymous):

Then just solve for y.

OpenStudy (anonymous):

24

OpenStudy (anonymous):

@Rav

OpenStudy (anonymous):

Did you solve for y? Show your work.